# HILCodec: High Fidelity and Lightweight Neural Audio Codec

### v1

@@ Abstract
The recent advancement of end-to-end neural audio codecs enables compressing audio at very low bitrates while reconstructing the output audio with high fidelity. Nonetheless, such improvements often come at the cost of increased model complexity. In this paper, we identify and address the problems of existing neural audio codecs. We show that the performance of Wave-U-Net does not increase consistently as the network depth increases. We analyze the root cause of such a phenomenon and suggest a variance-constrained design. Also, we reveal various distortions in previous waveform domain discriminators and propose a novel distortion-free discriminator. The resulting model, HILCodec, is a real-time streaming audio codec that demonstrates state-of-the-art quality across various bitrates and audio types.
@@ I Introduction
An audio codec is a system that compresses and decompresses audio data. It comprises an encoder which analyzes the input audio, a quantizer, and a decoder that reconstructs the output audio. Audio codecs have several objectives. First, they aim to compress the input using the minimum number of bits. Second, they strive to preserve the perceptual quality of the output, making it as similar as possible to the original input. Third, they seek to maintain a low delay for streaming applications. Lastly, they are designed to have low computational complexity so that they can be used on a variety of devices, including mobile and embedded systems.
@@
Traditional audio codecs leverage the expertise of speech processing, audio signal processing, and psychoacoustics to achieve their objectives. For speech coding, based on the source-filter model of speech production [2], linear predictive coding (LPC) [3] is widely adopted because of its high compression rate. Instead of compressing the waveform directly, LPC represents the input speech as an autoregressive model and quantizes the model’s coefficients. These coefficients are then dequantized and used to synthesize the output speech. However, LPC is not good at capturing high frequency components. Therefore, for other types of audios, such as music, transform coding is employed [4, 8, 11]. This method first transforms the input waveform into another domain, such as subbands [13], wavelet coefficients [15], or a time-frequency spectrogram [14], to obtain less correlated components. It then allocates available bits so that quantization noise for each component is psychoacoustically well-distributed [12]. Finally the decoder applies an inverse transform to obtain the output waveform.
@@
Recently, the remarkable advancements of machine learning have spurred active research into audio coding based on deep neural networks (DNNs). One research direction involves using a DNN as a post-processing enhancement module for existing codecs to suppress coding artifacts at low bitrates [30, 31, 32]. In some studies, a DNN decoder is trained on top of a traditional human-designed encoder. The input for such neural decoder can be the output representations of an encoder after quantization [33, 35], or the raw bitstream itself [34]. Another promising area is an end-to-end neural audio codec [36, 37, 38, 39, 42, 40, 41], where instead of relying on manually designed encoders and quantizers, the entire encoder-quantizer-decoder framework is trained in a data-driven way. Leveraging the power of generative adversarial networks (GAN) [43], end-to-end neural audio codecs have demonstrated state-of-the-art audio quality under low bitrate conditions. On the other hand, these improvements in audio quality are often attributed to an increase in network size, leading to substantial computational complexity. Furthermore, some of these codecs are not streamable. While there exist neural audio codecs that can be processed in real time streaming manner, it results in compromised performance. Given the growing demand and potential for neural audio codecs in real-world services, the development of a codec that is lightweight, streamable, and maintains high fidelity presents a timely and valuable opportunity.
In this paper, we challenge the common practice of simply augmenting model complexity. Our main focus is on dissecting and addressing the issues discovered from the existing generators and discriminators which hinder model performance. Combined with a lightweight backbone network, we propose HILCodec, an end-to-end neural audio codec that satisfies the aforementioned requirements. It is capable of compressing diverse types of audio data at bitrates ranging from 1.5 kbps to 9.0 kbps, with a sampling rate of 24kHz. HILCodec can operate in real-time on a single thread of a CPU while demonstrating comparable or superior audio quality to other traditional and neural audio codecs.
For the generator, we construct a low-complexity version of the widely-used Wave-U-Net architecture [16] using depthwise-separable convolutions [51]. We incorporate $L_{2}$-normalization and multi-resolution spectrogram inputs, which enhance the model’s performance with a minimal increase in computational complexity. These modifications serves us a room to increase the width and depth of the network.
However, we identify an issue of the Wave-U-Net architecture where an increase in network depth does not always leads to an improvement in audio quality and may even result in quality degradation. This problem has not been addressed explicitly in previous studies on audio coding. Through theoretical and empirical analysis, we deduce that the discrepancy is due to the exponential increase in variance of activations during the forward pass. To mitigate this, we propose a variance-constrained design, which effectively scales performance in line with network depth.
In addition, we show that existing waveform domain discriminators often operate on distorted signals. As a solution, we introduce a distortion-free discriminator that utilizes multiple filter banks to enhance perceptual quality. When combined with a time-frequency domain discriminator, it surpasses the performance of conventional discriminators in GAN-based audio codecs in terms of subjective quality. These innovative approaches significantly contribute to the overall effectiveness of our proposed HILCodec.
@@
  • We identify a problem associated with the Wave-U-Net generator: increasing network depth does not necessarily lead to improved performance. We address this by adopting a variance-constrained design.
@@ II-A Neural Vocoder
A vocoder is a system that synthesizes human voice from latent representations. The study on vocoders originated from [1] and has since been central to digital speech coding. Recently, neural vocoders have brought significant attention due to their applicability to text-to-speech (TTS) systems. A neural vocoder synthesizes speech from a given mel-spectrogram, while a separate neural network generates the mel-spectrogram from text. Various neural vocoders have been proposed based on deep generative models, including autoregressive models [18, 20, 19], flow [21], diffusion [25, 26], and others. Among these, vocoders based on the GAN framework have been popular due to their high-quality output and fast inference time. MelGAN [22] introduced the Multi-Scale Discriminator (MSD) which operates on average-pooled waveform and a feature matching loss to stabilize adversarial training. Parallel WaveGAN [23] introduced a multi-resolution short-time Fourier transform (STFT) loss, which outperforms the single-resolution counterpart. HiFi-GAN [24] proposed a well-designed generator and Multi-Period Discriminator (MPD) which operates on downsampled waveform. Avocodo [27] aimed to reduce aliasing in the generator and waveform domain discriminators by introducing the Cooperative Multi-Band Discriminator (CoMBD) and Subband Discriminator (SBD). Some models used time-frequency domain discriminators [28, 29]. Due to the similarity of the tasks, developments in the field of neural vocoders have been exploited for various neural audio codecs. In this work, our approach HILCodec also utilizes GAN-based training techniques from the line of this trend, while analysing signal distortion incurred by existing discriminators and proposing a novel filter-bank based distortion-free discriminator. This approach aligns with a method proposed in [27]. Nonetheless, compared to previous works, our discriminator shows superior generalization capabilities and can handle distortions around the transition bands.
@@ II-B End-to-end Neural Audio Codec
Soundstream [38] is a GAN-based end-to-end streaming neural audio codec trained on general audio. By utilizing residual vector quantizer [5, 6], it facilitates the training of a single model that supports multiple bitrates. EnCodec [39] introduced a novel loss balancer to stabilize training. Audiodec [40] proposed an efficient GAN training scheme and modified the model architecture of HiFi-GAN. Despite decent quality for speech coding, it requires 4 threads of a powerful workstation CPU for real-time operation. HiFi-Codec [41] and Descript Audio Codec [42] also adopted the generator and discriminators from HiFi-GAN, each showing state-of-the-art performance for clean speech and general audio compression. However, they are not streamable and their computational complexity is high. In this paper, we introduce a model architecture that is streamable and has low computational complexity. Furthermore, we pinpoint and address a previously unexplored problem within the waveform domain generator, proposing a solution that enhances audio quality beyond that of prior models.
@@ III Generator
III-A Model Architecture
HILCodec is an end-to-end neural audio codec that leverages the time-domain fully convolutional Wave-U-Net. Following [38], HilCodec is composed of an encoder, a residual vector quantizer (RVQ), and a decoder. Given a raw waveform input, the encoder generates a feature vector for every 320 samples. The decoder reconstructs the waveform from a sequence of quantized features. The overall model architecture is illustrated in Fig. 1.
III-A1 Encoder-Decoder
Similar to [17], the encoder is composed of a convolution layer with $C_{enc}$ output channels followed by 4 convolution blocks and a quantization block. Each encoder block contains a spectrogram block (see Section III-A4) and $N_{enc}-1$ residual blocks, followed by a strided convolution for downsampling. The strides are set to 2, 4, 5, and 8. The quantization block is composed of a convolution layer with $D$ output channels, an $L_{2}$-normalization layer (see Section III-A3), and an RVQ with codebook vectors of $D$ dimensions.
The decoder mirrors the encoder’s structure, replacing strided convolutions with transposed convolutions for upsampling, and substituting the quantization block with a dequantization block. Each decoder block contains $N_{dec}$ residual blocks. The final convolution layer has $C_{dec}$ input channels. An ELU activation function [50] is inserted before every convolution layer, except the first one of the encoder. A hyperbolic tangent activation is inserted after the last convolution layer of the decoder to ensure that the final output waveform lies within the range of $-1$ and $1$.
@@ III-A1 Encoder-Decoder
To ensure streamability, all convolution layers are causal. They are designed as depthwise separable convolutions [51] which help to maintain performance while reducing computation. Additionally, we apply weight normalization [52] to further enhance the model’s performance.
III-A2 Residual Vector Quantizer
@@ III-A3 $L_{2}$-normalization
Since the advent of VQVAE [55], there has been a continous exploration for superior vector quantization. One such method is the $L_{2}$-normalization of the encoder outputs and codebook vectors, which has been proven to improve the codebook utilization and the final output quality of a generative model [56, 42]. Our experiments also corroborate its beneficial impact on audio quality (see Section VI-B). However, unlike [56] and [42], we do not directly apply $L_{2}$-normalization to the codebook vectors. As we employ RVQ, we naturally apply the $L_{2}$-normalization only to the encoder output before quantization. This method bolsters the training stability and enhances the audio quality of the model.
@@
Figure 2: Comparison of the average channel variance according to network depth. On the x-axis, each number represents the network depth of the encoder and the decoder, and the terms “In”, “Q”, and “Out” denote the input waveform, the output of the residual vector quantizer, and the output waveform, respectively. The logaritmic y-scale on the left corresponds to the baseline model, while the linear y-scale on the right corresponds to the variance constrained model. The solid lines represent the residual blocks.
III-A4 Spectrogram Block
@@
Traditional audio codecs exploit human-designed feature extractors for compression. In contrast, an end-to-end neural codec learns a feature extractor in a data-driven manner without requiring any expertise on audio coding. While a larger model can learn better features, it comes at the cost of increased computational complexity. To circumvent this, we extract additional features akin to those used in transform coding and feed them into the encoder network. Specifically, we insert a spectrogram block at the beginning of each encoder block. We provide a log spectrogram whose time resolution and frequency size match the time resolution and channel size of the encoder block’s input respectively. These multi-resolution spectrogram inputs assist the encoder with a limited parameter size in learning richer features. Importantly, we preserve the streaming nature of our model by using a causal STFT to obtain the spectrogram.
III-B Variance-Constrained Design
III-B1 Problems of Wave-U-Net
@@
The Wave-U-Net architecture, which forms the backbone of the HilCodec generator, is a widely adopted model for end-to-end neural audio codecs [38, 39, 42, 41, 40]. However, we observed that the audio quality does not improve consistently as the depth of the model increases (see Fig. 5). This observation aligns with the results reported in [39], where the objective score for audio quality did not increase even after the number of residual blocks in encoder and decoder was tripled. To identify the reason, we concentrate on the signal propagation within the model’s forward pass.
@@
Fig. 2 illustrates the signal propagation plot [62] of HILCodec right after initialization. Note that the y-axis of the baseline model is on a logarithmic scale ranging from $10^{-3}$ to $10^{3}$. In the encoder, the variance increases exponentially until it reaches the final layer, the $L_{2}$-normalization layer. In the decoder, the exponential growth of the variance persists until the end. The reasons and downsides for such a phenomenon can be well understood in the literature on deep normalized networks.
@@
In the field of computer vision, the combination of BatchNorm [45] and skip connection [47] showed great performance gain and enabled very deep networks to be stably trained. Theoretical analyses have been conducted on the signal propagation of these networks to explain their success. Consider a network solely composed of residual blocks. The $n$-th residual block takes the form $\mathbf{x}_{n+1}=\mathbf{x}_{n}+f_{n}(\mathbf{x}_{n})$, where $\mathbf{x}_{n}$ is the input and $\mathbf{x}_{n+1}$ is the output. We refer to $f_{n}$ as the residual branch. Right after initialization, given that the input and the model weights are independent, $\mathbf{Var}(\mathbf{x}_{n+1})=\mathbf{Var}(\mathbf{x}_{n})+\mathbf{Var}(f_{n}(\mathbf{x}_{n}))$. Reference [61] showed that if $f_{n}$ is a network without any normalization layer, $\mathbf{Var}(f(\mathbf{x}_{n}))\approx\mathbf{Var}(\mathbf{x}_{n})$, leading to $\mathbf{Var}(\mathbf{x}_{n+1})\approx 2\mathbf{Var}(\mathbf{x}_{n})$. This implies that the variance increases exponentially according to the network depth $n$. On the other hand, if $f_{n}$ is a normalized network with appropriate initialization, $\mathbf{Var}(f_{n}(\mathbf{x}_{n}))\approx 1$, resulting in $\mathbf{Var}(\mathbf{x}_{n+1})\approx n+1$. This linear growth is beneficial to the model’s signal propagation -- it implies that the output of the residual branch $f_{n}(\mathbf{x}_{n})$ contributes only a $\frac{1}{n+1}$ fraction to $\mathbf{x}_{n+1}$. As a result, the output of each residual block is dominated by a skip connection, leading to stable training even for very deep networks. On the other hand, in un-normalized networks, both a residual branch and skip connection contribute equally to the output of each residual block, hindering signal propagation.
@@
Since the Wave-U-Net architecture doesn’t contain any normalization layer, it experiences an exponential growth in variance as a function of network depth, resulting in performance degradation as illustrated in Fig. 5. Also, due to the high dynamic range of the baseline model’s signal propagation, mixed-precision training [54] fails when the number of residual blocks in each encoder/decoder block exceeds two. Prior works have indirectly addressed this issue by increasing the network’s width instead of its depth. For instance, HiFi-GAN employs multi-receptive field fusion modules, where the outputs from multiple parallel residual blocks are aggregated to yield the final output, instead of sequentially connecting each block. Furthermore, AudioDec extends this structure by explicitly broadening the network width using grouped convolutions.
@@
Our goal is to directly solve the problem of exponential variance growth, rather than merely circumventing it. The most straightforward solution is to incorporate a normalization layer into every residual block. Nevertheless, our preliminary experiments with various normalization layers, including batch normalization and layer normalization layers [46], resulted in even poorer performance. We can only conjecture the reasons -- perhaps the batch size was insufficient due to hardware constraints, or the normalization layers we used were ill-suited for the waveform domain. To address this, instead of relying on normalization layers, we propose a variance-constrained design which facilitates proper signal propagation. This encompasses the residual block design, input and output normalization, and zero initialization of residual branches. As shown in Fig. 5, after applying our methods, the audio quality improves according to the network depth.
III-B2 Variance-Constrained Residual Block

### v2

# HILCodec: High-Fidelity and Lightweight Neural Audio Codec

@@ Abstract
The recent advancement of end-to-end neural audio codecs enables compressing audio at very low bitrates while reconstructing the output audio with high fidelity. Nonetheless, such improvements often come at the cost of increased model complexity. In this paper, we identify and address the problems of existing neural audio codecs. We show that the performance of the SEANet-based codec does not increase consistently as the network depth increases. We analyze the root cause of such a phenomenon and suggest a variance-constrained design. Also, we reveal various distortions in previous waveform domain discriminators and propose a novel distortion-free discriminator. The resulting model, HILCodec, is a real-time streaming audio codec that demonstrates state-of-the-art quality across various bitrates and audio types.
@@ I Introduction
An audio codec is a system that compresses and decompresses audio data. It comprises an encoder which analyzes the input audio, a quantizer, and a decoder that reconstructs the output audio. Audio codecs have several objectives [73]. First, they aim to compress the input using the minimum number of bits. Second, they strive to preserve the perceptual quality of the output, making it as similar as possible to the original input. Third, they seek to maintain a low algorithmic delay for streaming applications. Lastly, they are designed to have low computational complexity so that they can be used on a variety of devices, including mobile and embedded systems.
@@
Traditional audio codecs leverage the expertise of speech processing, audio signal processing, and psychoacoustics to achieve their objectives. For speech coding, based on the source-filter model of speech production [2], linear predictive coding (LPC) [3] is widely adopted because of its high compression rate. Instead of compressing the waveform directly, LPC represents the input speech as an autoregressive model and quantizes the model’s coefficients. These coefficients are then dequantized and used to synthesize the output speech. However, the limitation of LPC in capturing high-frequency components leads to the use of transform coding [4, 8, 11, 77, 78], particularly when preserving high-frequency components is crucial. This method first transforms the input waveform into another domain, such as sub-bands [13], wavelet coefficients [15], or a time-frequency spectrogram [14], to obtain less correlated components. It then allocates available bits so that quantization noise for each component is psychoacoustically well-distributed [12]. Finally, the decoder applies an inverse transform to obtain the output waveform.
@@
Recently, the remarkable advancements of machine learning have spurred active research into audio coding based on deep neural networks (DNNs). One research direction involves using a DNN as a post-processing enhancement module for existing codecs to suppress coding artifacts at low bitrates [30, 31, 32]. In some studies, a DNN decoder is trained on top of a traditional human-designed encoder. The input for such neural decoder can be the output representations of an encoder after quantization [33, 35], or the raw bitstream itself [34]. Another promising area is an end-to-end neural audio codec [36, 37, 38, 39, 42, 40, 41], where instead of relying on manually designed encoders and quantizers, the entire encoder-quantizer-decoder framework is trained in a data-driven way. Leveraging the power of generative adversarial networks (GAN) [43], end-to-end neural audio codecs have demonstrated state-of-the-art audio quality under low bitrate conditions. On the other hand, these improvements in audio quality are often attributed to an increase in network size, leading to substantial computational complexity. Furthermore, some of these codecs are not streamable. While there exist neural audio codecs that can be processed in real time streaming manner, it results in compromised performance.
In this paper, we aim to develop a neural audio codec model with the following properties:
  • Streaming: When an input audio is divided into short chunks and provided in a streaming manner, the model must be able to encode and decode each chunk sequentially.
  • Lightweight: The number of model parameters should be small.
  • Low-complexity: The number of multiply-accumulate (MAC) opeartions should be low.
  • High-fidelity: The codec’s output should be perceptually similar to the input across various bitrates and audio types.
To fulfill streaming, lightweight, and low-complexity capabilities, we designed a lightweight version of the widely used SEANet architecture [16] (Section III). For high fidelity, our strategy is three-fold. First, we incorporate $L_{2}$-normalization (Section III-C) and multi-resolution spectrogram inputs (Section III-D), which enhance the model’s performance and training stability with a minimal increase in computational complexity. Second, we identify an issue with the SEANet architecture where an increase in network depth does not always lead to an improvement in audio quality and may even result in quality degradation (Section III-E1). This problem has not been addressed explicitly in previous studies on audio coding. Through theoretical and empirical analysis, we deduce that this discrepancy is due to the exponential increase in variance of activations during the forward pass. To mitigate this, we propose a variance-constrained design, which effectively scales performance in line with network depth (Section III-E2 to Section III-E5). Lastly, we show that existing waveform domain discriminators often distort input signals before feeding them into neural networks (Section IV-A1). As a solution, we introduce a distortion-free discriminator that utilizes multiple filter banks to enhance perceptual quality (Section IV-A2). When combined with a time-frequency domain discriminator, it surpasses the performance of conventional discriminators in GAN-based audio codecs in terms of subjective quality. These innovative approaches significantly contribute to the overall effectiveness of our proposed model, HILCodec.
@@
  • We identify a problem associated with the SEANet generator: increasing network depth does not necessarily lead to improved performance. We address this by adopting a variance-constrained design.
@@ II-A Neural Vocoder
A vocoder is a system that synthesizes human voice from latent representations. The study on vocoders originated from [1] and has since been central to digital speech coding. Recently, neural vocoders have brought significant attention due to their applicability to text-to-speech (TTS) systems. A neural vocoder synthesizes speech from a given mel-spectrogram, while a separate neural network generates the mel-spectrogram from text. Various neural vocoders have been proposed based on deep generative models, including autoregressive models [18, 20, 19], flow [21], diffusion [25, 26], and others. Among these, vocoders based on the GAN framework have been popular due to their high-quality output and fast inference time. MelGAN [22] introduced the Multi-Scale Discriminator (MSD) which operates on a waveform after average pooling, and a feature matching loss to stabilize adversarial training. Parallel WaveGAN [23] introduced a multi-resolution short-time Fourier transform (STFT) loss, which outperforms the single-resolution counterpart. HiFi-GAN [24] proposed a well-designed generator and the Multi-Period Discriminator (MPD) which operates on downsampled waveform. Avocodo [27] aimed to reduce aliasing in the generator and waveform domain discriminators by introducing the Cooperative Multi-Band Discriminator (CoMBD) and the Sub-Band Discriminator (SBD). Some models used time-frequency domain discriminators [28, 29]. Due to the similarity of the tasks, developments in the field of neural vocoders have been exploited for various neural audio codecs. In this work, we also utilize GAN-based training techniques from the line of this trend, while analyzing signal distortion incurred by existing discriminators and proposing a novel filter bank-based distortion-free discriminator. This approach aligns with a method proposed in [27]. Nonetheless, compared to previous works, our discriminator can handle distortions around the transition bands and shows superior audio quality.
@@ II-B End-to-end Neural Audio Codec
Soundstream [38] is a GAN-based end-to-end streaming neural audio codec trained on general audio. By utilizing residual vector quantizer [5, 6], it facilitates the training of a single model that supports multiple bitrates. EnCodec [39] introduced a novel loss balancer to stabilize training. AudioDec [40] proposed an efficient GAN training scheme and modified the model architecture of HiFi-GAN. Despite decent quality for speech coding, it requires 4 threads of a powerful workstation CPU for real-time operation. HiFi-Codec [41] and Descript Audio Codec [42] also adopted the generator and discriminators from HiFi-GAN, each showing state-of-the-art performance for clean speech and general audio compression. However, they are not streamable and their computational complexity is high. FreqCodec [74] proposed a lightweight frequency domain model with low computational complexity, showing exceptional performance in objective evaluations. However, it did not undergo a subjective evaluation. Penguins [75] and LightCodec [76] employed low-complexity neural coding modules, each demonstrating subjective quality superior to EnCodec. Nonetheless, their evaluations were limited to a speech corpus. Additionally, these models are not publicly available, and their capabilities for streaming remain unspecified. In this paper, we introduce a model architecture that is streamable, lightweight, and has low computational complexity. Furthermore, we pinpoint and address a previously unexplored problem within the waveform domain generator, proposing a solution that enhances audio quality beyond that of prior models over various audio types.
@@ III Generator
HILCodec is an end-to-end neural audio codec that leverages the time-domain fully convolutional layers. The overall model architecture is illustrated in Fig. 1. Following [38], HilCodec is composed of an encoder, a residual vector quantizer (RVQ), and a decoder (Section III-A, III-B). Given a raw waveform input, the encoder generates a feature vector for every 320 samples. The decoder reconstructs the waveform from a sequence of quantized features. It is designed in favor of four desired properties: streaming, lightweight, low-complexity, and high-fidelity. To ensure the streaming property, all convolutions are causal [17]. To achieve lightweight and low-complexity properties, we utilize depthwise separable convolutions [51]. To achieve high-fidelity without significantly increasing the model complexity, we introduce $L_{2}$-normalization (Section III-C), spectrogram blocks (Section III-D), and a variance-constrained design (Section III-E). In the following subsections, we provide a detailed explanation of each part of the model.
III-A Encoder-Decoder
Similar to [16], the encoder is composed of a convolutional layer with $C_{enc}$ output channels followed by 4 encoder blocks and a quantization block. Each encoder block contains a spectrogram block (Section III-D) and $N_{enc}-1$ residual blocks, followed by a strided convolution for downsampling. The strides are set to 2, 4, 5, and 8. The quantization block is composed of a spectrogram block, two convolutions, an $L_{2}$-normalization layer (Section III-C), and an RVQ with codebook vectors of $128$ dimensions.
The decoder mirrors the encoder’s structure, replacing strided convolutions with transposed convolutions for upsampling, and substituting the quantization block with a dequantization block. The dequantization block is composed of a dequantization layer, a pointwise convolution, and a depthwise convolution. Each decoder block contains $N_{dec}$ residual blocks. The final convolution has $C_{dec}$ input channels. An ELU activation function [50] is inserted before every convolution, except the first one of the encoder. A hyperbolic tangent activation is inserted after the last convolution of the decoder to ensure that the final output waveform lies within the range of $-1$ and $1$. Following [39], we apply weight normalization [52] to every convolution to further enhance the model’s performance.
@@ III-A Encoder-Decoder
III-B Residual Vector Quantizer
@@ III-C $L_{2}$-normalization
Since the advent of VQVAE [55], there has been a continuous exploration for superior vector quantization. One such method is the $L_{2}$-normalization of the encoder outputs and codebook vectors, which has been proven to improve the codebook utilization and the final output quality of a generative model [56, 42]. In our experiments, although $L_{2}$-normalization did not consistently improve audio quality across all audio types, we applied it to our model because it stabilizes mixed-precision training [54] (Section VI-B1). Unlike [56] and [42], we do not directly apply $L_{2}$-normalization to the codebook vectors. As we employ RVQ, we naturally apply the $L_{2}$-normalization only to the encoder output before quantization.
@@
Figure 2: Comparison of the average channel variance after initialization according to network depth. On the x-axis, each number represents the network depth of the encoder and the decoder, and the terms “In”, “Q”, and “Out” denote the input waveform, the output of the residual vector quantizer, and the output waveform, respectively. The logaritmic y-scale on the left corresponds to the baseline model, while the linear y-scale on the right corresponds to the variance constrained model. The solid lines represent the residual blocks.
III-D Spectrogram Block
@@
Traditional audio codecs exploit human-designed feature extractors for compression. In contrast, an end-to-end neural codec learns a feature extractor in a data-driven manner without requiring any expertise on audio coding. While a larger model can learn better features, it comes at the cost of increased computational complexity. To circumvent this, we extract additional features akin to those used in transform coding and feed them into the encoder network. Specifically, we insert a spectrogram block at the beginning of each encoder block and quantization block. As shown in Fig. 1, a spectrogram block is composed of a log-magnitude spectrogram extractor, a normalization layer (Section III-E4), a pointwise convolution, and a skip connection. We provide log-magnitude spectrograms with hop sizes of $\{1,2,8,40,320\}$ and Fourier transform sizes of $\{64,128,256,512,1024\}$. The hop sizes are set so that the time resolution of each spectrogram matches that of the input of each encoder block and quantization block. These multi-resolution spectrograms assist the encoder with a limited parameter size in learning richer features. Importantly, we preserve the streaming nature of our model by using a causal STFT to obtain the spectrogram.
III-E Variance-Constrained Design
For deep neural networks, proper signal propagation in the early stages of training is particularly important. However, we have identified that the existing model architecture exhibits undesired signal propagation immediately after initialization (Section III-E1). As a solution, we propose a variance-constrained design (Sections III-E2 to Section III-E5). This approach enhances audio quality without increasing model complexity (Section VI-B1).
III-E1 Variance Explosion Problem of SEANet
@@
The SEANet architecture, which forms the backbone of the HilCodec generator, is a widely adopted model for end-to-end neural audio codecs [38, 39, 42, 41, 40]. However, we observed that the audio quality does not improve consistently as the depth of the model increases (Fig. 6). This observation aligns with the results reported in [39, Table A.3], where the objective score for audio quality did not increase even after the number of residual blocks in the encoder and decoder was tripled. If the dataset size is limited, the performance of a DNN model may not improve as the model size increases. However, since our model is trained on a large audio corpus (Section V-B), it is natural to expect that the performance scales with the model size. To investigate this discrepancy, we concentrate on the signal propagation within the model’s forward pass.
@@
Fig. 2 illustrates the signal propagation plot [62] of HILCodec right after initialization. Note that the y-axis of the baseline model is on a logarithmic scale ranging from $10^{-3}$ to $10^{3}$. In the encoder, the variance increases exponentially until it reaches the final layer, the $L_{2}$-normalization layer. In the decoder, the exponential growth of the variance persists until the end. The reasons and downsides for such a variance explosion phenomenon can be well understood in the literature on deep normalized networks [61, 62, 70].
@@
In the field of computer vision, the combination of batch normalization [45] and skip connection [47] showed great performance gain and enabled very deep networks to be stably trained. Theoretical analyses have been conducted on the signal propagation of these networks to explain their success. Consider a network solely composed of residual blocks. The $n$-th residual block takes the form $\mathbf{x}_{n+1}=\mathbf{x}_{n}+f_{n}(\mathbf{x}_{n})$, where $\mathbf{x}_{n}$ is the input and $\mathbf{x}_{n+1}$ is the output. We refer to $f_{n}$ as the residual branch. Right after initialization, given that the input and the model weights are independent, $\mathbf{Var}(\mathbf{x}_{n+1})=\mathbf{Var}(\mathbf{x}_{n})+\mathbf{Var}(f_{n}(\mathbf{x}_{n}))$. Reference [61] showed that if $f_{n}$ is a properly initialized network without any normalization layer, $\mathbf{Var}(f(\mathbf{x}_{n}))\approx\mathbf{Var}(\mathbf{x}_{n})$, leading to $\mathbf{Var}(\mathbf{x}_{n+1})\approx 2\mathbf{Var}(\mathbf{x}_{n})$. This implies that the variance increases exponentially according to the network depth $n$. On the other hand, if $f_{n}$ is a normalized network with appropriate initialization, $\mathbf{Var}(f_{n}(\mathbf{x}_{n}))\approx 1$, resulting in $\mathbf{Var}(\mathbf{x}_{n+1})\approx n+1$. This linear growth is beneficial to the model’s signal propagation -- it implies that the output of the residual branch $f_{n}(\mathbf{x}_{n})$ contributes only a $\frac{1}{n+1}$ fraction to $\mathbf{x}_{n+1}$. As a result, the output of each residual block is dominated by a skip connection, leading to stable training even for very deep networks. On the other hand, in unnormalized networks, both the residual branch and skip connection contribute equally to the output of each residual block, hindering signal propagation at the beginning of training.
@@
Back to our model, since the SEANet architecture doesn’t contain any normalization layer, it experiences an exponential growth in variance as a function of network depth, resulting in performance degradation as illustrated in Fig. 6. Also, due to the high dynamic range of the baseline model’s signal propagation, mixed-precision training fails when the number of residual blocks in each encoder/decoder block exceeds two. Prior works have indirectly addressed this issue by increasing the network’s width instead of its depth. For instance, HiFi-GAN employs multi-receptive field fusion modules, where the outputs from multiple parallel residual blocks are aggregated to yield the final output, instead of sequentially connecting each block. Furthermore, AudioDec extends this structure by explicitly broadening the network width using grouped convolutions.
@@
Our goal is to directly solve the problem of exponential variance growth that occurs immediately after initialization, rather than merely circumventing it. The most straightforward solution is to incorporate a normalization layer into every residual block. Nevertheless, our preliminary experiments with various normalization layers, including batch normalization and layer normalization [46], resulted in even poorer performance. We can only conjecture the reasons -- perhaps the batch size was insufficient due to hardware constraints, or the normalization layers we used were ill-suited for the waveform domain. To address this, instead of relying on normalization layers, we propose a variance-constrained design (VCD) which facilitates proper signal propagation. This encompasses the residual block design, input and output normalization, and zero initialization of residual branches. As shown in Fig. 6, after applying our methods, the audio quality improves according to the network depth. Note that the proposed design is not constrained to the SEANet architecture -- it can be applied to various models that have skip connections.
III-E2 Variance-Constrained Residual Block